# Identity Preserving 3D Head Stylization with Multiview Score Distillation

### v1

@@ 4 Experiments
Baselines. We include various domain adaptation methods for our comparisons and adopt all of them for the PanoHead generator. StyleCLIP [40] trains a $\mathcal{W}^{+}$ mapper network with the CLIP loss. StyleGAN-NADA [13] employs the CLIP objective to train adaptively-selected generator layers rather than modifying $\mathcal{W}^{+}$. StyleGANFusion [51] utilizes SDS for domain adaptation and a directional regularizer via the frozen generator for more stable distillation. DiffusionGAN3D [29] further improves the regularization of [51] by employing a relative distance loss. After tuning the generators, for inference, we invert images by optimizing $\mathcal{W}^{+}$ to reconstruct the input image with the original PanoHead generator. The optimized $\mathcal{W}^{+}$ is then fed to these different domain-adapted generators, and rendering from various views is outputted. Details of the baseline implementations are detailed in Supplementary.
@@
Specifically, in FID and CLIP, we construct two different edited image distributions, wherein one, the images are edited with a Stable Diffusion pipeline in the 2D image domain (ground-truth distribution), and in the other, the images are edited via the domain-adapted 3D-aware generators. We provide the example images from the ground-truth distribution in Supplementary. We calculate FID between those two image distributions and CLIP scores between paired images.
In ID and $\Delta\mathcal{D}$, we use ground truth unedited images and calculate the scores between the ground-truth original images and stylized images.
Results. We provide quantitative and qualitative results in Tabs. 1, 1 and 4, respectively. The quantitative results demonstrate that our method outperforms others in terms of stylization quality, as measured by FID and CLIP scores, as well as in identity preservation, as evaluated by ArcFace-based ID similarity and multi-view depth difference $\Delta\mathcal{D}$ across nearly all prompts. These numerical findings are consistent with the visual assessment. As shown in Figs. 1 and 4, StyleCLIP and StyleGAN-NADA produce stylizations of lower quality, with noticeable artifacts in the renderings. While StyleGANFusion and DiffusionGAN3D are capable of producing stylized images, they significantly compromise identity preservation. For instance, in the Joker example in Fig. 1, both methods generate identical facial features for two different identities. Similarly, for the sketch prompt, the identities are substantially altered. Competing methods also remove accessories such as eyeglasses. In contrast, our method effectively preserves the identity of the input image while delivering high-quality stylizations.
@@
StyleCLIP [40] 118.2 0.77 0.52 0.022 97.6 0.75 0.38 0.031 248.8 0.63 0.42 0.039 103.7 0.75 0.54 0.026 181.6 0.63 0.61 0.012
StyleGAN-NADA [13] 81.1 0.81 0.61 0.013 116.3 0.81 0.50 0.011 212.9 0.75 0.45 0.012 99.7 0.71 0.51 0.034 154.5 0.76 0.36 0.039
StyleGANFusion [51] 168.0 0.76 0.60 0.003 119.3 0.80 0.41 0.007 203.9 0.70 0.41 0.012 99.0 0.74 0.53 0.012 85.5 0.83 0.47 0.007
DiffusionGAN3D [29] 189.3 0.82 0.46 0.024 110.7 0.86 0.47 0.009 132.3 0.81 0.68 0.002 159.4 0.71 0.48 0.016 166.2 0.82 0.43 0.009
@@
Table 1: Quantitative scores with competitive methods. The first and second best results are shown in bold and underlined.
@@
Figure 5: Mesh and rendering visualizations of ours and competing methods for the Pixar prompt. The input consists of a single image, and we display the inverted and rendered results from PanoHead from three different viewpoints.
In Fig. 5, we present the mesh predictions for the Pixar-style stylizations produced by our method and the competing approaches. The first row shows reconstructions of the input image from different views, both in RGB image space and from the mesh perspective. These results demonstrate that our method achieves significantly better identity preservation both in terms of facial features and distinctive accessories, such as eyeglasses. Unlike other methods, which completely remove or distort the eyeglasses, our approach maintains them with high fidelity.
Figure 6: Percentage of users preferences are presented for different prompts. Users overwhelmingly select our method when compared to the other approaches.
We then conduct a user study with 25 participants to evaluate the quality of 3D stylization and identity preservation across different methods. Participants are shown images generated by five different models: StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, and our own approach. For each image, they are asked to select which output best balances stylization and identity preservation. The methods are presented in random order for each image to minimize bias. Results of this user study are shown in Fig. 6. The data indicate that our method is consistently preferred across all the prompts tested, with participants overwhelmingly selecting it as the best for both stylization and identity preservation when compared to the other approaches.
Ablation Study. We begin our ablation study with the DiffusionGAN3D [29] baseline and build upon it by incorporating various improvements. [29] uses SDS loss for generator optimization, along with a relative distance loss. This relative distance loss aims to preserve the distance between two images generated by the original generator, which serves as the ground truth. The fine-tuned generator is then optimized to maintain this same distance between the corresponding generated samples.
@@
Figure 7: Qualitative Ablation Study I. LD and other regularizers like grid and mirror scores help keep the original hair color, skin tone, and geometry while stylizing.
As shown in Fig. 7, while their method successfully achieves stylization, the results fail to preserve the characteristics of the input images. Next, we replace the text-based diffusion model that generates the SDS loss with the depth-based ControlNet [62], which helps retain more characteristics from the input image, though its effectiveness remains limited. In the next setting, we apply negative log-likelihood distillation (LD) alongside the depth-based ControlNet model. This results in less saturation, and we observe improved preservation of the input identity. We also show our final results demonstrating the improved color and geometry alignment with the input images that include the improvements from grid and mirror gradients in the last column. These visuals are consistent with the quantitative improvements presented in Tab. 2.
@@
[29] 123.4 - 0.76 - 0.53 -
[29] + [62] 108.3 15.1 0.76 0 0.58 0.05
LD + [62] 93.0 15.3 0.77 0.01 0.76 0.18
LD + [62] + Weighted rank 91.4 1.6 0.76 0.01 0.77 0.01
+ Grid after SR 86.7 4.7 0.79 0.03 0.70 0.07
+ Grid before SR 85.7 1.0 0.79 0 0.72 0.02
+ Mirror gradients 82.3 3.4 0.80 0.01 0.72 0
Table 2: Quantitative ablation study. $\Delta$ represents the difference between the results of the current row and the row above it. We use the prompts ”sketch”, ”statue”, and ”werewolf”, and present their averaged results.
Input LD + [62] weighted-rank Grid after SR Grid before SR Grid + mirror
@@
Results with other prompts. We present additional results with various prompts in Fig. 9. These prompts include local stylization, such as adding a mustache or editing hair to pink. For these experiments, we do not employ attribute-based masking for score gradients (for example, we do not mask the face when editing the hair). Despite this, our method successfully achieves both localized edits and more global transformations, such as modifying the character to resemble an elf under a single framework.
@@
Metallic face paint Refer to caption
Pink hair Refer to caption
@@ 5 Conclusion
We introduced a novel approach to 3D head stylization that improves identity preservation and stylization quality. By utilizing the PanoHead model for 360-degree image synthesis and incorporating negative log-likelihood distillation (LD), mirror and grid score gradients, and score rank weighing, our method overcomes the limitations of previous approaches that struggled with identity retention. Experimental results show significant qualitative and quantitative improvements, advancing the state of 3D head stylization. Our work also provides valuable insights into effective distillation processes between diffusion models and GANs, emphasizing the importance of identity preservation in stylization tasks.

### v2

@@ 4 Experiments
Baselines. We include various domain adaptation methods for our comparisons and adopt all of them for the PanoHead generator. StyleCLIP [40] trains a $\mathcal{W}^{+}$ mapper network with the CLIP loss. StyleGAN-NADA [14] employs the CLIP objective to train adaptively-selected generator layers rather than modifying $\mathcal{W}^{+}$. StyleGANFusion [50] utilizes SDS and a directional regularizer via the frozen generator for more stable distillation. DiffusionGAN3D [29] further improves the regularization of [50] by employing a relative distance loss. After tuning the generators, for inference, we invert images by optimizing $\mathcal{W}^{+}$ to reconstruct the input image with the original PanoHead generator. The optimized $\mathcal{W}^{+}$ is then fed to these different domain-adapted generators, and rendering from various views is outputted. We also include 2D-3D editing methods such as InstantID [53], InstructPix2Pix [8] and InstructNerf2Nerf [16] for a broader comparison. Details are given in Supplementary.
@@
Specifically, in FID and CLIP, we construct two different edited image distributions, wherein one, the images are edited with a Stable Diffusion pipeline in the 2D image domain (ground-truth distribution), and in the other, the images are edited via the domain-adapted 3D-aware generators. We provide the example images from the ground-truth distribution in Supplementary. We calculate FID between those two image distributions and CLIP scores between paired images. In ID and $\Delta\mathcal{D}$, we use ground truth unedited images and calculate the scores between the ground-truth original images and stylized images.
Results. We provide quantitative and qualitative results in Tabs. 1, 1, 6 and 8, respectively. The quantitative results demonstrate that our method outperforms others in terms of stylization quality, as measured by FID and CLIP scores, as well as in identity preservation, as evaluated by ID similarity and $\Delta\mathcal{D}$ across nearly all prompts. These numerical findings are consistent with the visual assessment. As shown in Figs. 1 and 6, StyleCLIP and StyleGAN-NADA produce stylizations of lower quality with noticeable artifacts. While StyleGANFusion and DiffusionGAN3D are capable of producing stylized images, they significantly compromise identity preservation. For instance, in the Joker example in Fig. 1, both methods generate identical facial features for two different identities. Similarly, for the sketch prompt, the identities are substantially altered. Competing methods such as InstantID and IP2P also remove accessories such as eyeglasses and alter identity heavily on different views (Fig. 8). In contrast, ours effectively preserves the identity while delivering high-quality stylizations.
@@
2D InstructPix2Pix 144.4 0.82 0.40 0.024 116.8 0.90 0.44 0.023 178.1 0.82 0.28 0.013 89.5 0.65 0.24 0.002 60.9 0.91 0.39 0.011
InstantID 160.8 0.68 0.59 0.050 170.3 0.69 0.47 0.066 183.3 0.58 0.53 0.065 162.7 0.65 0.59 0.047 160.4 0.71 0.56 0.038
3D domain apt. StyleCLIP 118.2 0.77 0.52 0.022 97.6 0.75 0.38 0.031 248.8 0.63 0.42 0.039 103.7 0.75 0.54 0.026 181.6 0.63 0.61 0.012
StyleGAN-NADA 81.1 0.81 0.61 0.013 116.3 0.81 0.50 0.011 212.9 0.75 0.45 0.012 99.7 0.71 0.51 0.034 154.5 0.76 0.36 0.039
StyleGANFusion 168.0 0.76 0.60 0.003 119.3 0.80 0.41 0.007 203.9 0.70 0.41 0.012 99.0 0.74 0.53 0.012 85.5 0.83 0.47 0.007
DiffusionGAN3D 189.3 0.82 0.46 0.024 110.7 0.86 0.47 0.009 132.3 0.81 0.68 0.002 159.4 0.71 0.48 0.016 166.2 0.82 0.43 0.009
@@
Table 1: Quantitative scores with competitive domain adaptation and 2D-3D editing methods.
@@
Figure 7: Mesh and rendering visualizations of ours and competing methods for the Pixar prompt.
In Fig. 7, we present the mesh predictions for the Pixar-style stylizations produced by our method and the competing approaches. The first row shows reconstructions of the input image from different views, in RGB and mesh. The results demonstrate that our method achieves significantly better identity preservation both in terms of facial features and distinctive accessories, such as eyeglasses, unlike other methods which completely remove or distort them.
Input Pixar Joker Werewolf Sketch
Ours
Refer to caption Refer to caption Refer to caption Refer to caption
IN2N
Refer to caption Refer to caption Refer to caption Refer to caption
IP2P
Refer to caption Refer to caption Refer to caption Refer to caption
InstantID
Ours
Refer to caption Refer to caption Refer to caption Refer to caption
IN2N
Refer to caption Refer to caption Refer to caption Refer to caption
IP2P
Refer to caption Refer to caption Refer to caption Refer to caption
InstantID
Figure 8: Additional comparisons with other 2D-3D editing methods. 2D methods are run with the same seed for different views.
Ablation Study. We begin our ablation study with the DiffusionGAN3D [29] baseline and build upon it by incorporating various improvements. [29] uses SDS for generator optimization, along with a relative distance loss. This relative distance loss aims to preserve the distance between two images generated by the original generator. The fine-tuned generator is then optimized to maintain this same distance between the corresponding generated samples.
@@
Figure 9: Qualitative Ablation Study I. LD and other regularizers like grid and mirror scores help keep the original attributes.
As shown in Fig. 9, while their method successfully achieves stylization, the results fail to preserve the characteristics of the input images. Next, we attach depth-based ControlNet [63] to the text-based diffusion model, which helps retain more characteristics from the input image, though its effectiveness remains limited. In the next setting, we apply negative log-likelihood distillation (LD) alongside the depth-based ControlNet. This results in less saturation, and we observe improved identity preservation. We also show our final results demonstrating the improved color and geometry alignment with the input images that include the improvements from grid and mirror gradients in the last column. These visuals are consistent with the quantitative improvements presented in Tab. 2.
@@
Runtime. Our method requires $\sim$3h of generator tuning per prompt on a single A100 GPU. After tuning, inference on any head image takes only $\sim$0.5s. In comparison, [16] requires $\sim$1h to fit NeRF to a single identity and +3h for stylizing the scene. [53, 8] take 50 iterations through denoiser, taking $\sim$10-15s. [29, 50] also tune their generators, and their training & inference times are similar to ours.
@@
[29] 147.19 - 0.74 - 0.35 -
[29] + [63] 140.20 6.99 0.75 0.01 0.34 0.01
LD + [63] 116.88 23.32 0.81 0.06 0.46 0.12
LD + [63] + Weighted rank 116.09 0.79 0.82 0.01 0.48 0.02
+ Grid after SR 99.53 16.56 0.85 0.03 0.49 0.01
+ Grid before SR 81.11 18.42 0.85 0.00 0.51 0.02
+ Mirror gradients 84.17 3.06 0.86 0.01 0.52 0.01
Table 2: Quantitative ablation study. $\Delta$ represents the difference between the results of the current row and the row above. Our approach improves stylization with ID retention.
Input LD + [63] $\mathbf{\Sigma}$-weigh Grid after SR Grid before SR Grid + mirror
@@
Results with other prompts. We present additional results with various prompts in Fig. 11. These prompts include local stylization, such as adding a mustache or editing hair to pink. For these experiments, we do not employ attribute-based masking for score gradients (for example, we do not mask the face when editing the hair). Despite this, our method successfully achieves such transformations.
@@ 5 Conclusion
We introduced a novel approach to 3D head stylization that improves identity preservation and stylization quality. By utilizing the PanoHead model for 360-degree image synthesis and incorporating negative log-likelihood distillation (LD), mirror and grid score gradients, and score rank weighing, our method overcomes the limitations of previous approaches that struggled with identity retention. Experimental results show significant qualitative and quantitative improvements, advancing the state of 3D head stylization. Our work also provides valuable insights into effective distillation between diffusion models and GANs, emphasizing the importance of ID preservation in stylization.
Input & Joker Input & Sketch
Refer to caption Refer to caption
Figure 12: Limitations in preserving earrings and hair accessories.
Limitations and future work. A limitation, similar to other methods, is that stylization may not fully preserve earrings and hair accessories, as illustrated in Fig. 12. Another one is the output resolution, which is bounded by the utilized generative model. However, they can be mitigated by utilizing a heavier model. For the future work, we will extend our ID preserving approach and findings about domain adaptation to class-agnostic, large reconstruction models. We believe utilizing LD instead of SDS, score rank-weighing, and grid denoising for such pre-trained models will unlock new research directions in efficient and controllable domain adaptation and stylization.
[…]
Supplementary Material
@@ Appendix B Implementation details
Rank weighing on score tensors. Fig. 14 illustrates how an SVD‐based approach can decompose a stylized portrait into coarse and fine components, and then reconstruct it at different “rank” levels (k). This progressive refinement underlines how SVD can serve as a powerful control mechanism for score distillation, letting the user dial in how many spatial “frequencies” of the style are included.
Input $\mathbf{\Sigma}$-weighted $k=1$ $k=2$ $k=3$ $k=4$
Caricature Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Charcoal Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Statue Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Oil Painting Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Pixar Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Werewolf Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Zombie Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption Refer to caption
Figure 14: Demonstration of our SVD‐based stylization across multiple face styles. The Input column shows the original images. $\Sigma$‐weighted uses all top singular values with decreasing weights, preserving both coarse and fine features. Columns k=1 through k=4 depict rank‐k approximations; as k increases, more high‐frequency details are retained, resulting in sharper, more faithful stylizations.
Appendix C User study
Figure 15: Percentage of user preferences. Users overwhelmingly favor ours compared to other domain adaptation methods.
We conduct a user study with 25 participants to evaluate the quality of 3D stylization and identity preservation across different methods. Participants are shown images generated by five different models: StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, and our own approach. For each image, they are asked to select which output best balances stylization and identity preservation. The methods are presented in random order for each image to minimize bias. Results of this user study are shown in Fig. 15. The data indicate our method is consistently preferred across all the prompts tested, with participants overwhelmingly selecting it as the best for both stylization and identity preservation compared to others.
Appendix D Additional results
Pixar Joker Werewolf Sketch Statue
2D InstructPix2Pix  0.1461  0.1164  0.1427  0.0790  0.0900
InstantID  0.0897  0.1185  0.1055  0.1218  0.1290
3D StyleCLIP  0.1045  0.0958  0.1962  0.0878  0.1658
StyleGAN-NADA  0.0459  0.1380  0.2617  0.0890  0.1480
StyleGANFusion  0.1668  0.1904  0.1387  0.1168  0.1212
DiffusionGAN3D  0.1566  0.0977  0.0922  0.1216  0.2442
Ours  0.0326  0.0713  0.0856  0.0742  0.1122
Table 3: KID scores on the test set.
Tab. 3 reveals the KID scores on the same test set used in the main paper. Our method outperforms all baselines in KID across domains, with the exception of the Statue domain. Notably, in the Sketch stylization setting, while InstructPix2Pix reports a slightly better FID, our method achieves superior KID scores.
Figs. 16, 18, 17, 21, 19 and 20 visualizes the outputs of methods for different prompts in 360-degrees.
Figure 16: Joker edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.
Figure 17: Pixar edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.
Figure 18: Sketch edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.
Figure 19: Werewolf edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.
Figure 20: Zombie edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.
Figure 21: Statue edits. From top to bottom: input, StyleCLIP, StyleGAN-NADA, StyleGANFusion, DiffusionGAN3D, ours.